OpenStudy (anonymous):

Graph the following linear system to determine the number of solutions. y equals negative three x plus six, six x plus two y equals twelve zero 1 2 infinitely many

OpenStudy (anonymous):

plzz asap i need help it for a grade

OpenStudy (goten77):

y= -3x + 6 y= 6x +2 = 12 not sure what your saying....???

OpenStudy (anonymous):

y = -3x + 6 6x + 2y = 12 lets rearrange the first equation 3x + y = 6 now we have: 3x + y = 6 (2) = 6x + 2y = 12 6x + 2y = 12 there is infinite solutions